OpenStudy (anonymous):

Find three consecutive even integers such that the sum of the first integer and three times the third integer is 64 more than the second integer

OpenStudy (anonymous):

If you assume n is the smallest of the even integers, how could you write the next two interns of n, since they are 2 and 4 units away, respectively?

OpenStudy (zzr0ck3r):

hint: letting n be the first even integer, then the next is n+2, what would the next be?

OpenStudy (anonymous):

Supoose the smallest integer is 18, the using zzr0ck3r's expression the next integer would be 18 +2 = 20. What is the third of the integers in terms of n?

OpenStudy (anonymous):

I get the point thanks

OpenStudy (anonymous):

Can you set-up the equation?

OpenStudy (anonymous):

n+3(n+4)=n+66

OpenStudy (anonymous):

Yes!!! you are right. I must be getting punchy!

OpenStudy (anonymous):

Now you are set. Solve for n, then complete the sequence.

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

You are welcome.